Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get a 280lb mixture that is 12% protein?
Found 3 solutions by checkley77, stanbon, mananth:
Answer by checkley77(12844)

.14x+.07(280-x)=.12*280
.14x+19.6-.07x=33.6
.07x=33.6-19.6
.07x=14
x=14/.07
x=200 lbs. of soybeans is used.
280-200=80 lbs. of cornmeal is used.
Proof:
.14*200+.07*80=.12*280
28+5.6=33.6
33.6=33.6

Answer by stanbon(75887)

Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get a 280lb mixture that is 12% protein?
--------------------
Equation:
0.14x + 0.07(280-x) = 0.12*280
---
Multiply thru by 100 to get:
14x + 7*280 - 7x = 12*280
7x = 5*280
x = 5*40
x = 200 lb (amt of soybeal meal needed for the mixture)
---
280-x = 80 lb (amt of cornmeal needed for the mixture)
===========================================================
Cheers,
Stan H.

Answer by mananth(16946)

Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get a 280lb mixture that is 12% protein?
soyabean - 14% protien
cornmeal - 7% protien
280 -12 %
Let soyabean required be x lbs
cornmeal will be 280-x lbs
0.14x+0.07(280-x)=0.12*280
0.14x+19.60-0.07x=33.60
0.07x = 33.60 -19.60
0.07x=14.00
x= 14/ 0.07
x= 200 pounds of soyabean
280-200 = 80 pounds of corn meal is required
Ananth